OpenStudy (anonymous):

In the function f(x) = 5(x2 − 4x + ____) + 15, what number belongs in the blank to complete the square? Numerical Answers Expected!

OpenStudy (anonymous):

4

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

you have -4x which is made up of -2x and -2x now -2 times -2 = 4 thats why...

OpenStudy (anonymous):

Solve 2x2 + 12x − 14 = 0 by completing the square. 2(x + 6)2 = 26; x = −6 + square root 13, x = −6 − square root 13 2(x + 6)2 = 20; x = −6 + square root 20, x = −6 − square root 20 2(x + 3)2 = 32; x = 1, x = −7 2(x + 3)2 = 32; x = 13, x = 19

OpenStudy (anonymous):

x2 - 4x +4 = (x-2)(x-2) = (x-2)^2 this is completing the square

OpenStudy (anonymous):

can you help me with that question too? please

OpenStudy (anonymous):

so which answers would it be?

OpenStudy (anonymous):

can u give it a try first :)

OpenStudy (anonymous):

i want u to learn something here

OpenStudy (anonymous):

umm well i don't think its a

OpenStudy (anonymous):

first simplify it mate

OpenStudy (anonymous):

2 ( x^2 + 6x - 7)

OpenStudy (anonymous):

now u have +6x which is made up of +3x and +3x, so u have 3 and 3 , when u multiply them u get 9

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do you make the -7 to a +9 ???

OpenStudy (anonymous):

16

OpenStudy (anonymous):

well done

OpenStudy (anonymous):

so we get this, pay attention and tell me why did I put 32 outside, ok?

OpenStudy (anonymous):

2( x^2 + 6x -7 + 16) - 32 = 0

OpenStudy (anonymous):

2 (x^2 + 6x + 9) = 32

OpenStudy (anonymous):

following me ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

divide both side by 2 and factorize

OpenStudy (anonymous):

(x+3)^2 = 16 what is x ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

one value for x is 1 the other is -7 why -7 though? remember that when u Square a negative number u get a positive one, right ? so the answer is... 2(x + 3)2 = 32; x = 1, x = −7 now why did u get 5 ??

OpenStudy (anonymous):

i didn't pay attention the "2" was an exponent

OpenStudy (anonymous):

Two sets of quadratic expressions are shown below in various forms. Line 1: x2 + 3x + 2 (x + 1)(x + 2) (x + 1.5)2 − 0.25 Line 2: x2 + 5x + 6 (x + 2)(x + 3) (x + 2.5)2 + 6.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (anonymous):

ok, so are you happy with this and do u understand it, or want to ask something about it? dont be shy to ask, ok?

OpenStudy (anonymous):

no but i do have another question

OpenStudy (anonymous):

in Line 1: x2 + 3x + 2 = (x + 1)(x + 2) and in Line 2: x2 + 5x + 6 = (x + 2)(x + 3) so you need to confirm the end bits of the lines if they are equivalent to the first 2, can u do that by yourself?

OpenStudy (anonymous):

Two sets of quadratic expressions are shown below in various forms. Line 1: x2 + 3x + 2 (x + 1)(x + 2) (x + 1.5)2 − 0.25 Line 2: x2 + 5x + 6 (x + 2)(x + 3) (x + 2.5)2 + 6.25 Which line contains three equivalent expressions? Line 1 only Line 2 only Both Line 1 and Line 2 Neither Line 1 nor Line 2

OpenStudy (anonymous):

is it line 1?

OpenStudy (anonymous):

it could be both, you'll know once u expand the end bits

OpenStudy (anonymous):

I'll help u with the first end bit, but u do the second one ok?

OpenStudy (anonymous):

how do i do that then

OpenStudy (anonymous):

(x + 1.5)2 − 0.25 (x+1.5)(x+1.5) - 0.25 x^2 + 1.5x + 1.5x + 2.25 - 0.25 x^2 + 3x + 2 now is this equivalent to x2 + 3x + 2 and (x + 1)(x + 2) ??

OpenStudy (anonymous):

can u do the second one and show me?

OpenStudy (anonymous):

no not really lol

OpenStudy (anonymous):

lol..why

OpenStudy (anonymous):

i dont knwo how

OpenStudy (anonymous):

(x + 2.5)2 + 6.25 (x + 2.5)(x + 2.5) + 6.25 x^2 + 2.5x + 2.5x + 6.25 + 6.25 x^2 +5x + 12.5 and this is NOT equivaent to the previous x2 + 5x + 6 and (x + 2)(x + 3)

OpenStudy (anonymous):

so ur earlier guess of " Line 1 only" was correct but it was only a guess, u need to practice more on factorizing and simplifying if u want to be able to do maths with ease :)

OpenStudy (anonymous):

okay thank you i have one last question

OpenStudy (anonymous):

The cost to produce a product is modeled by the function f(x) = 5x2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. 5(x − 7)2 + 13; The minimum cost to produce the product is $13. 5(x − 7)2 + 13; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $258.

OpenStudy (anonymous):

The cost to produce a product is modeled by the function f(x) = 5x2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. 5(x − 7)2 + 13; The minimum cost to produce the product is $13. 5(x − 7)2 + 13; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $7. 5(x − 7)2 + 258; The minimum cost to produce the product is $258.

OpenStudy (anonymous):

5(x − 7)2 + 13; The minimum cost to produce the product is $13. 5(x − 7)2 + 13; The minimum cost to produce the product is $7. one of the above

OpenStudy (anonymous):

is it the second one?

OpenStudy (anonymous):

I think its the first one

OpenStudy (anonymous):

i could be wrong tho, hold on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

damn...am confused, please ask someone else, do u mind?...sorry...